OpenStudy (anonymous):

Simplify (6ab^3c)(-abc^2).

OpenStudy (anonymous):

-6a^2b^4c^3

OpenStudy (anonymous):

-6a^2b^4c^3

OpenStudy (anonymous):

see the terms in first bracket can be written as 6 * a * b^3 * c the term in second bracket can be written as -1 * a * b * c^2 so their product = 6 * a * b^3 * c * -1 * a * b * c^2 = 6 * -1 * a * a * b^3 * b * c * c^2 = -6 * a^2 * b^4 * c^3 = -6a^2b^4c^3